Hi i need help with ny practice test that my teacher gave me and is big package and im doing " Percent Word Problems"
1.what percent of 126 is 22 ?
2. what percent of 137.4 is 96?
3.79.9 is 90% of what ?
4.73% of what is 156.4 ?

Answer by rfer(16322)

1)
22/126=0.175=17.5%
2)
96/137.4=0.70=70%
3)
79.9/0.90=88.78
4)
156.4/0.0473=3306.55